OpenStudy (mikezack123):

A car is traveling at 30 m/s with a Kinetic Energy of 900 kJ what is its mass

OpenStudy (anonymous):

do you know the formula for kinetic energy?

OpenStudy (mikezack123):

No and I'm on a time limit, please help.

OpenStudy (anonymous):

The formula is ... kinetic energy= (mass/2)*speed^2

OpenStudy (anonymous):

\(kinetic~energy=\frac{ mass }{ 2 }*speed^2\)

OpenStudy (anonymous):

Now plug in the numbers you know, which are the SPEED and the KINETIC ENERGY

OpenStudy (anonymous):

After you do this, solve it as an equation

OpenStudy (mikezack123):

I don't get it

OpenStudy (mikezack123):

Oh wait I have to solve for m?

OpenStudy (mikezack123):

Omg help I'm running out of time

OpenStudy (mikezack123):

Please

OpenStudy (anonymous):

And yes, you have to solve for M

OpenStudy (mikezack123):

?

OpenStudy (anonymous):

Solve for M, you know the speed and kinetic energy,

OpenStudy (anonymous):

Solve this... \[900=\frac{ m }{ 2 }*900\]

OpenStudy (mikezack123):

M=2

OpenStudy (mikezack123):

So the mass of the car is 2 kg?